OpenStudy (anonymous):

Please help Q1. 4) 2R=1800 INCORRECT MY R=9K AND rm=9k......... Q 6. , V0=5V, T=0.5ms, Rw=3Ω, Lw=1mH and RL=2Ω. 3)? 4)? I am using the correct formuka but not the right answers...... help

OpenStudy (anonymous):

3 to tu ne karadia tha yar???

OpenStudy (anonymous):

Q1 4) 18k

OpenStudy (anonymous):

sure????

OpenStudy (anonymous):

no doubt in dat

OpenStudy (anonymous):

Q6 c)1.828 d)0.15

OpenStudy (anonymous):

@Abhaypratap 18k is wrong for Q1 part 4.. that was my last chance :(

OpenStudy (anonymous):

hmmm i thnk i wrote the small 'k' thats why

OpenStudy (anonymous):

oops in question they mentioned ans to b written in terms of kohms i should hav reminded u dat sorry :(

OpenStudy (anonymous):

shall I write 18K in capital?

OpenStudy (anonymous):

Abhaypratap

OpenStudy (anonymous):

just write 18

OpenStudy (anonymous):

thanx alot I ve got 93% overall

OpenStudy (anonymous):

My Pleasure :) congrats

OpenStudy (anonymous):

what Q6 point A and B

OpenStudy (anonymous):

a) 0.2 b) 0.2

OpenStudy (anonymous):

as per d question in d post

OpenStudy (anonymous):

plz guys elps me on Q1 1 2 and 3 that what i have as value The elements in the circuit have the following values: R=5kΩ and rm=15kΩ

OpenStudy (anonymous):

a)ia=2*i1*R/(2*R+rm) and vo=ia*rm b)iA =(-i2)*R/(2*R+rm) and V0=ia*rm c) part a) + part b) d) R+R

OpenStudy (anonymous):

here are the answeres 0.2 0.2 1.828 0.15 dont forget best response